# On weight and variance uncertainty in neural networks for regression tasks

### v1

@@ 5 Variational Bayes Regression Network with Variance Uncertainty
In this section, we suggest to use a variational posterior distribution over weights, biases, and also the variance of the likelihood function. Our proposal can be considered as a generalization of the previous model introduced in section 4. In our approach, we assume that the posterior parameters are represented by $\theta=(W,S)$, where $W$ encompasses the weights and biases of the NN, and $S$ denotes a parameter related to the variance of the likelihood function. Additionally, we define the variational parameters as $\eta=(\sigma_{w}^{2},\mu_{w},\sigma_{L}^{2},\mu_{L})$, where $\sigma_{w}^{2}$ and $\mu_{w}$ pertain to the variational distribution over $W$, and $\sigma_{L}^{2}$ and $\mu_{L}$ are associated to variational distribution over $S$.
@@
$y_{i}=\phi\left(X_{i};W\right)+\varepsilon_{i}={\hat{y}}_{i}+\varepsilon_{i},\ \ \ \ \ \ \varepsilon_{i}\sim N(0,\ \log\left(1+\exp(S\right))I_{q})$
that gives
$y_{i}|x_{i}\sim N({\hat{y}}_{i},\ \log\left(1+\exp(S\right))I_{q}).$
@@
$L(W,S|x,y)$ $=\prod_{i=1}^{n}\frac{1}{\sqrt{2\pi\log\left(1+\exp(S\right))}}\exp\left\{-\frac{1}{2\log\left(1+\exp(S\right))}(y_{i}-\phi(x_{i};W))^{T}(y_{i}-\phi(x_{i};W))\right\}.$
Based on [4], we assume that the variational posterior of all parameters is a diagonal Gaussian distribution, i.e., $W|\mu_{w},\sigma_{w}^{2}\sim N(\mu_{w},\sigma_{w}^{2}I_{|W|})$, where $|W|$ denotes the number of weights and biases in the NN, and $\sigma_{w}=\log\left(1+\exp(\rho_{w})\right)>0$. Also let
@@
such that
$\sigma_{L}^{2}=\log\left(1+\exp(\rho_{L})\right).$
Further, assume $\varepsilon_{w}\sim\mathcal{N}(0,I_{|W|})$ and $\varepsilon_{L}\sim\mathcal{N}(0,1)$. Then, the posterior samples of $W$ and $S$ are given by
$W=\mu_{w}+\varepsilon_{w}\odot\log(1+\exp(\rho_{w}))$
$S=\mu_{L}+\varepsilon_{L}\log(1+\exp(\rho_{L})),$
where $\odot$ denotes element-wise multiplication. Here, $\mu_{w}$ and $\rho_{w}$ are the variational parameters associated with $W$, and $\mu_{L}$ and $\rho_{L}$ are the variational parameters associated with $S$.
@@
  1. Sample $\varepsilon_{w}\sim\mathcal{N}(0,I_{|W|})$] and $\varepsilon_{L}\sim\mathcal{N}(0,1)$.
  2. Compute posterior sample:
$W=\mu_{w}+\varepsilon_{w}\odot\log(1+\exp(\rho_{w})).$
$S=\mu_{L}+\varepsilon_{L}\log(1+\exp(\rho_{L}))$
@@ 3. Define the objective function:
$f(W,\mu_{w},\sigma_{w}^{2})=\log q(W|\mu_{w},\sigma_{w}^{2})-\log p(W,S)p(D|W,S).$
$f(S,\mu_{L},\sigma_{L}^{2})=\log q(S|\mu_{L},\sigma_{L}^{2})-\log p(W,S)p(D|W,S).$
  4. Calculate the gradient of $f(W,\mu_{w},\sigma_{w}^{2})$ with respect to $\mu_{w}$ and $\rho_{w}$ and the gradient of $f(S,\mu_{L},\sigma_{L}^{2})$ with respect to $\mu_{L}$ and $\rho_{L}$.
$\nabla_{\mu_{w}}f(W,\mu_{w},\sigma_{w}^{2})=\frac{\partial f(W,\mu_{w},\sigma_{w}^{2})}{\partial\mu_{w}},\nabla_{\rho_{w}}f(W,\mu_{w},\sigma_{w}^{2})=\frac{\partial f(W,\mu_{w},\sigma_{w}^{2})}{\partial\rho_{w}},$
$\nabla_{\mu_{L}}f(S,\mu_{L},\sigma_{L}^{2})=\frac{\partial f(S,\mu_{L},\sigma_{L}^{2})}{\partial\mu_{L}},\nabla_{\rho_{L}}f(S,\mu_{L},\sigma_{L}^{2})=\frac{\partial f(S,\mu_{L},\sigma_{L}^{2})}{\partial\rho_{L}}.$
  5. Update the variational parameters.
$\mu_{w}\leftarrow\mu_{w}-\gamma_{w}\,\nabla_{\mu_{w}}f(W,\mu_{w},\sigma_{w}^{2}),\rho_{w}\leftarrow\rho_{w}-\gamma_{w}\,\nabla_{\rho_{w}}f(W,\mu_{w},\sigma_{w}^{2})$
$\mu_{L}\leftarrow\mu_{L}-\gamma_{l}\,\nabla_{\mu_{L}}f(S,\mu_{L},\sigma_{L}^{2}),\rho_{L}\leftarrow\rho_{L}-\gamma_{l}\,\nabla_{\rho_{L}}f(S,\mu_{L},\sigma_{L}^{2})$
  6. Repeat from step 1 until the stopping condition is satisfied.
@@ 6 Experimental Results
    $p(W)=\frac{1}{\sqrt{2\pi\sigma_{p}^{2}}}\exp\left(-\frac{1}{2\sigma_{p}^{2}}vec(W)^{\top}vec(W)\right),$
    where $\sigma_{p}^{2}$ denotes the variance of the prior, and $\theta$ represents the vector of parameters.
@@
    $p(W)=\prod_{j}\left[z_{j}\mathcal{N}(W_{j};0,\sigma_{1}^{2})+(1-z_{j})\mathcal{N}(W_{j};0,\sigma_{2}^{2})\right],$
@@ 6.1 Nonlinear function estimation
$y=x+2\sin\left(2\pi(x+\epsilon)\right)+2\sin\left(4\pi(x+\epsilon)\right)+\epsilon,$
where $\epsilon\sim N(0,0.02)$. To model the out-of-sample unforeseen risks, the values of $x$ for the train and test samples are generated from uniform distributions with supports $[-0.1,0.6]$ and $[-0.25,0.85]$, respectively. The train and test sets contain 800 and 200 samples, respectively.
@@
The prior $p(W)=\mathcal{N}(W;0,\sigma_{p}^{2}I_{|W|})$ is used for the VBNET-FIXED, while and the prior $p(W,S)=\mathcal{N}(W;0,\sigma_{pw}^{2}I_{|W|})\mathcal{N}(S;0,\sigma_{ps}^{2})$ is considered for the VBNET-SVAR. Four competitive models are considered for the comparison study; VBNET-SVAR, VBNET-FIXED, the standard NN (NNET), and the Generalized Additive Model (GAM [9]). A total number of 10 replications is performed in this simulation study.
@@ 6.2 Riboflavin dataset
Riboflavin is a dataset with a small number of samples but many features. This makes it a suitable dataset for Bayesian models. Dataset of riboflavin production by Bacillus subtilis [5] contains n = 71 observations of p = 4088 predictors (gene expressions) and a one-dimensional response (riboflavin production). From 71 samples, 56 randomly drawn samples are considered as the train set, and the remaining 25 samples are considered as the test set. We have repeated this sampling 10 times to evaluate the performance of the competitor models on this data set. All network was constructed using two hidden layers with 128 and 64 neurons, respectively.
Two scenarios are examined for fitting the NN on the train set. In the first scenario (PCA-BNN), the first 25 principal components of the 4088 genes are considered as the input of the neural network and all other competitive models. The second scenario (dropout-BNN) utilizes the feature selection using the dropout mechanism, imposed on the BNNs by the spike-and-slab prior. For the second scenario, the GAM model is replaced by the sparse-GAM model (GAMSEL) [6].
Figure 4: The result of samples 1-4 for the riboflavin data set, with PCA-BNN scenario.
Figure 5: The box plots of the MSPE for the riboflavin data set, with PCA-BNN scenario.
Figure 6: The result of samples 1-4 for the riboflavin data set, with dropout-BNN scenario.
Figure 7: The box plots of the MSPE for the riboflavin data set, with dropout-BNN scenario.
Figures 4 and 6 shows the predicted values for all competitive models, for the PCA-BNN and dropout-BNN scenarios, respectively, including 95% prediction intervals for the BNNs. Figures 5 and 7 presents box plots for the values of the MSPE for each model,for the PCA-BNN and dropout-BNN scenarios, respectively. From Figures 5 and 7, one can observe that the VBNET-SVAR model outperforms the other competitive models for out-of-sample prediction.

### v2

@@ 5 Variational Bayes Regression Network with Variance Uncertainty
In this section, we suggest to use a variational posterior distribution over weights, biases, and also the variance of the likelihood function. Our proposal can be considered as a generalization of the previous model introduced in section 4. In our approach, we assume that the posterior parameters are represented by $\boldsymbol{\theta}=(\mathbf{W},S)$, where $\mathbf{W}$ encompasses the weights and biases of the NN, and $S$ denotes a parameter related to the variance of the likelihood function. Additionally, we define the variational parameters as $\boldsymbol{\eta}=(\boldsymbol{\sigma}_{w}^{2},\boldsymbol{\mu}_{w},\sigma_{L}^{2},\mu_{L})$, where $\boldsymbol{\sigma}_{w}^{2}$ and $\boldsymbol{\mu}_{w}$ pertain to the variational distribution over $\mathbf{W}$, and $\sigma_{L}^{2}$ and $\mu_{L}$ are associated to variational distribution over $S$.
@@
$\mathbf{y}_{i}=\phi\left(\mathbf{x}_{i};W\right)+\varepsilon_{i}={\hat{\mathbf{y}}}_{i}+\varepsilon_{i},\ \ \ \ \ \ \varepsilon_{i}\sim N(0,\ {\color[rgb]{0,0,0}g(S)}\mathbf{I_{q}}),$
where $g:\mathbb{R}\rightarrow(0,\infty)$ is a strictly positive transformation function mapping the unconstrained posterior parameter $S$ to a valid variance value. In this work, we employ the softplus function, $g(S)=\log(1+\exp(S))$, to guarantee the positivity of the variance parameter. Consequently, the conditional distribution of $\mathbf{y}_{i}$ is given by:
$\mathbf{y}_{i}|\mathbf{x}_{i}\sim N({\hat{y}}_{i},{\color[rgb]{0,0,0}g(S)}\mathbf{I_{q}}),\quad{\color[rgb]{0,0,0}g(S)=\log(1+\exp\{S\})}.$
@@
$L(\mathbf{W},S|\mathbf{x},\mathbf{y})$ $=\prod_{i=1}^{n}\frac{1}{\sqrt{(2\pi)^{q}({\color[rgb]{0,0,0}g(S)})^{q}}}\exp\left\{-\frac{1}{2{\color[rgb]{0,0,0}g(S)}}(\mathbf{y}_{i}-\phi(\mathbf{x}_{i};\mathbf{W}))^{T}(\mathbf{y}_{i}-\phi(\mathbf{x}_{i};\mathbf{W}))\right\}.$
Based on [5] and utilizing the mean-field approximation $q(\mathbf{W},S)=q(\mathbf{W})q(S)$, we assume that the variational posterior of all parameters is a diagonal Gaussian distribution, i.e., $\mathbf{W}|\boldsymbol{\mu}_{w},\boldsymbol{\sigma}_{w}^{2}\sim N(\boldsymbol{\mu}_{w},\operatorname{diag}(\boldsymbol{\sigma}_{w}^{2}))$, where $|\mathbf{W}|$ denotes the number of weights and biases in the NN, and $\boldsymbol{\sigma}_{w}=\log\left(1+\exp(\boldsymbol{\rho}_{w})\right)>0$. Similarly, for the posterior parameter $S$, we assume
@@
where its standard deviation is parameterized as
$\color[rgb]{0,0,0}\sigma_{L}=\log\left(1+\exp(\rho_{L})\right).$}
Further, by assuming $\boldsymbol{\varepsilon}_{w}\sim\mathcal{N}(0,I_{|\mathbf{W}|})$ and $\varepsilon_{L}\sim\mathcal{N}(0,1)$, and applying the reparameterization trick, the posterior samples of $\mathbf{W}$ and $S$ are given by
$\mathbf{W}=\boldsymbol{\mu}_{w}+\boldsymbol{\varepsilon}_{w}\odot{\color[rgb]{0,0,0}\boldsymbol{\sigma}_{w}},$
$S=\mu_{L}+\varepsilon_{L}{\color[rgb]{0,0,0}\sigma_{L}},$
where $\odot$ denotes element-wise multiplication. Here, $\boldsymbol{\mu}_{w}$ and $\boldsymbol{\rho}_{w}$ are the variational parameters associated with $\mathbf{W}$, and $\mu_{L}$ and $\rho_{L}$ are the variational parameters associated with $S$.
Since $\boldsymbol{\sigma}_{w}$ and $\sigma_{L}$ are parameterized by $\boldsymbol{\rho}_{w}$ and $\rho_{L}$, respectively, the set of variational parameters to be optimized can be written as $\boldsymbol{\eta}=(\boldsymbol{\rho}_{w},\boldsymbol{\mu}_{w},\rho_{L},\mu_{L})$.
5.1 Objective function and gradient estimation
In order to minimize the negative $\rm ELBO$, with respect to the variational parameters $\boldsymbol{\eta}=(\boldsymbol{\rho}_{w},\boldsymbol{\mu}_{w},\rho_{L},\mu_{L})$, which involves intractable expectations, we employ Stochastic Gradient Descent (SGD). First, recall that the objective function (negative ELBO) for our proposed method can be written as:
$\mathcal{F}(\boldsymbol{\eta})=\mathbb{E}_{q(\mathbf{W},S|\boldsymbol{\eta})}[f(\mathbf{W},S,\boldsymbol{\eta})],$
where $q(\mathbf{W},S|\boldsymbol{\eta})=\mathcal{N}(\mathbf{W};\boldsymbol{\mu}_{w},\operatorname{diag}(\boldsymbol{\sigma}^{2}_{w}))\mathcal{N}(S;\mu_{L},\sigma^{2}_{L})$ is a diagonal Gaussian variational posterior over $\mathbf{W}$ and $S$, $p(\mathbf{W},S)=p(\mathbf{W})p(S)$ represents the factorized prior (typically Gaussian, acting as a regularizer), and $f(\mathbf{W},S,\boldsymbol{\eta})$ is the cost function defined as:
$f(\mathbf{W},S,\boldsymbol{\eta})=\log q(\mathbf{W}|\boldsymbol{\mu}_{w},\boldsymbol{\sigma}^{2}_{w})+\log q(S|\mu_{L},\sigma^{2}_{L})-\log p(\mathbf{W})-\log p(S)-\log L(\mathbf{W},S|\mathbf{x},\mathbf{y}).$
It is important to note that the KL divergence term is implicitly minimized via the term $\mathbb{E}_{q}[f(\mathbf{W},S,\boldsymbol{\eta})]$. We use a Monte Carlo approximation for this term rather than an analytic closed-form solution. This approach avoids the need for tedious derivation of the KL divergence for complex priors and allows for greater flexibility in the choice of prior distributions.
To minimize this objective, we compute the gradients of $\mathcal{F}(\boldsymbol{\eta})$ with respect to $\boldsymbol{\eta}$. Using standard properties of expectations (via the Dominated Convergence Theorem) and applying the reparameterization trick, the gradients can be expressed as expectations over the variable $\varepsilon$:
$\nabla_{\boldsymbol{\mu}_{w}}\mathcal{F}(\boldsymbol{\eta})=\mathbb{E}_{\boldsymbol{\varepsilon}_{w}\sim\mathcal{N}(0,I_{|\mathbf{W}|})}\left[\frac{\partial f}{\partial\mathbf{W}}+\frac{\partial f}{\partial\boldsymbol{\mu}_{w}}\right],$
$\nabla_{\boldsymbol{\rho}_{w}}\mathcal{F}(\boldsymbol{\eta})=\mathbb{E}_{\boldsymbol{\varepsilon}_{w}\sim\mathcal{N}(0,I_{|\mathbf{W}|})}\left[\frac{\partial f}{\partial\mathbf{W}}\odot\frac{\boldsymbol{\varepsilon}_{w}}{1+\exp\{-\boldsymbol{\rho}_{w}\}}+\frac{\partial f}{\partial\boldsymbol{\rho}_{w}}\right].$
Similarly, for the variational parameters $\mu_{L}$ and $\rho_{L}$, we have:
$\nabla_{\mu_{L}}\mathcal{F}(\boldsymbol{\eta})=\mathbb{E}_{\varepsilon_{L}\sim\mathcal{N}(0,1)}\left[\frac{\partial f}{\partial S}+\frac{\partial f}{\partial\mu_{L}}\right],$
$\nabla_{\rho_{L}}\mathcal{F}(\boldsymbol{\eta})=\mathbb{E}_{\varepsilon_{L}\sim\mathcal{N}(0,1)}\left[\frac{\partial f}{\partial S}\cdot\frac{\varepsilon_{L}}{1+\exp\{-\rho_{L}\}}+\frac{\partial f}{\partial\rho_{L}}\right].$
Since the expectations above are analytically intractable, we use a Monte Carlo estimator. Let $\boldsymbol{\varepsilon}^{1}_{w},...,\boldsymbol{\varepsilon}^{M}_{w}$ and $\varepsilon^{1}_{L},...,\varepsilon^{M}_{L}$ be i.i.d. samples from $\mathcal{N}(0,I_{|\mathbf{W}|})$ and $\mathcal{N}(0,1)$, respectively. By applying the reparameterization trick, let $\mathbf{W}_{j}=\boldsymbol{\mu}_{w}+\boldsymbol{\varepsilon}^{j}_{w}\odot\boldsymbol{\sigma}_{w}$ and $S_{j}=\mu_{L}+\varepsilon^{j}_{L}\sigma_{L}$ denote the posterior samples for j=1,,Mj=1,...,M. The gradient estimator for $\boldsymbol{\mu}_{w}$ is defined as:
$\widehat{\nabla}_{\boldsymbol{\mu}_{w}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{\partial f(\mathbf{W},S,\boldsymbol{\eta})}{\partial\mathbf{W}}+\frac{\partial f(\mathbf{W},S,\boldsymbol{\eta})}{\partial\boldsymbol{\mu}_{w}}\right]_{\mathbf{W}=\mathbf{W}_{j},S=S_{j}},$ (6)
which is an unbiased estimator. The estimators for $\boldsymbol{\rho}_{w},\mu_{L}$, and $\rho_{L}$ are defined analogously.
Note that the computational complexity of VBNET-SVAR is comparable to VBNET-FIXED, as only two additional scalar parameters have been introduced and is independent of size of the neural network.
@@ 5.1 Objective function and gradient estimation
1. Initialize the variational parameters $\boldsymbol{\eta}=(\boldsymbol{\mu}_{w},\boldsymbol{\rho}_{w},\mu_{L},\rho_{L})$ randomly.
2. Sample $\boldsymbol{\varepsilon}^{1}_{w},...,\boldsymbol{\varepsilon}^{M}_{w}\overset{\text{iid}}{\sim}\mathcal{N}(0,I_{|\mathbf{W}|})$ and $\varepsilon^{1}_{L},...,\varepsilon^{M}_{L}\overset{\text{iid}}{\sim}\mathcal{N}(0,1)$.
3. Compute posterior samples:
$\color[rgb]{0,0,0}\mathbf{W}_{j}=\boldsymbol{\mu}_{w}+\boldsymbol{\varepsilon}^{j}_{w}\odot\log(1+\exp(\boldsymbol{\rho}_{w})),\quad{\color[rgb]{0,0,0}S_{j}=\mu_{L}+\varepsilon^{j}_{L}\log(1+\exp(\rho_{L})),\quad j=1,...,M.}$}
@@ 4. Define the objective function:
$\color[rgb]{0,0,0}f(\mathbf{W},S,\boldsymbol{\eta})=\log q(\mathbf{W}|\boldsymbol{\mu}_{w},\boldsymbol{\sigma}_{w}^{2})+\log q(S|\mu_{L},\sigma_{L}^{2})-\log p(\mathbf{W})-\log p(S)-\log L(\mathbf{W},S|\mathbf{x},\mathbf{y}).$}
5. Calculate the gradients using equation (6):
$\color[rgb]{0,0,0}\widehat{\nabla}_{\boldsymbol{\mu}_{w}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{\partial f}{\partial\mathbf{W}}+\frac{\partial f}{\partial\boldsymbol{\mu}_{w}}\right]_{\mathbf{W}=\mathbf{W}_{j},S=S_{j}},$}
$\color[rgb]{0,0,0}\widehat{\nabla}_{\boldsymbol{\rho}_{w}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{\partial f}{\partial\mathbf{W}}\odot\frac{\boldsymbol{\varepsilon}^{j}_{w}}{1+\exp\{-\boldsymbol{\rho}_{w}\}}+\frac{\partial f}{\partial\boldsymbol{\rho}_{w}}\right]_{\mathbf{W}=\mathbf{W}_{j},S=S_{j}},$}
$\color[rgb]{0,0,0}\widehat{\nabla}_{\mu_{L}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{\partial f}{\partial S}+\frac{\partial f}{\partial\mu_{L}}\right]_{\textbf{W}=\textbf{W}_{j},S=S_{j}},$}
$\color[rgb]{0,0,0}\widehat{\nabla}_{\rho_{L}}=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{\partial f}{\partial S}\cdot\frac{\varepsilon^{j}_{L}}{1+\exp\{-\rho_{L}\}}+\frac{\partial f}{\partial\rho_{L}}\right]_{\textbf{W}=\textbf{W}_{j},S=S_{j}}.$}
6. Update the variational parameters:
$\color[rgb]{0,0,0}\boldsymbol{\mu}_{w}\leftarrow\boldsymbol{\mu}_{w}-\gamma_{w}\,\widehat{\nabla}_{\boldsymbol{\mu}_{w}},\quad\boldsymbol{\rho}_{w}\leftarrow\boldsymbol{\rho}_{w}-\gamma_{w}\,\widehat{\nabla}_{\boldsymbol{\rho}_{w}},$}
$\color[rgb]{0,0,0}\mu_{L}\leftarrow\mu_{L}-\gamma_{l}\,\widehat{\nabla}_{\mu_{L}},\quad\rho_{L}\leftarrow\rho_{L}-\gamma_{l}\,\widehat{\nabla}_{\rho_{L}}.$}
7. Repeat from step 2 until convergence.
@@ 6 Experimental Results
    $p(\mathbf{W})=\frac{1}{\sqrt{(2\pi\sigma_{p}^{2})^{|\mathbf{W}|}}}\exp\left(-\frac{1}{2\sigma_{p}^{2}}vec(\mathbf{W})^{\top}vec(\mathbf{W})\right),$
    where $\sigma_{p}^{2}$ denotes the variance of the prior, and W represents the vector of parameters.
@@
    $p(\mathbf{W})=\prod_{j}\left[z_{j}\mathcal{N}(W_{j};0,\sigma_{1}^{2})+(1-z_{j})\mathcal{N}(W_{j};0,\sigma_{2}^{2})\right],$
@@ 6.1 Nonlinear function estimation
$y=x+2\sin\left(2\pi(x+\varepsilon)\right)+2\sin\left(4\pi(x+\varepsilon)\right)+\varepsilon,$
where $\varepsilon\sim N(0,0.02)$. To model the out-of-sample unforeseen risks, the values of $x$ for the train and test samples are generated from uniform distributions with supports [0.1,0.6][-0.1,0.6] and [0.25,0.85][-0.25,0.85], respectively. The train and test sets contain 800 and 200 samples, respectively.
@@
The prior $p(\mathbf{W})=\mathcal{N}(\mathbf{W};0,\sigma_{p}^{2}I_{|\mathbf{W}|})$ is used for the VBNET-FIXED, while and the prior $p(\mathbf{W},S)=\mathcal{N}(\mathbf{W};0,\sigma_{pw}^{2}I_{|\mathbf{W}|})\mathcal{N}(S;0,\sigma_{ps}^{2})$ is considered for the VBNET-SVAR. Four competitive models are considered for the comparison study; VBNET-SVAR, VBNET-FIXED, the standard NN (NNET), and the Generalized Additive Model (GAM [10]). A total number of 10 replications is performed in this simulation study.
@@ 6.2 Riboflavin dataset: PCA-BNN Scenario
Riboflavin is a dataset with a small number of samples but many features. This makes it a suitable dataset for Bayesian models. Dataset of riboflavin production by Bacillus subtilis [6] contains n = 71 observations of p = 4088 predictors (gene expressions) and a one-dimensional response (riboflavin production). From 71 samples, 56 randomly drawn samples are considered as the train set, and the remaining 15 samples are considered as the test set. We have repeated this sampling 10 times to evaluate the performance of the competitor models on this data set.
In this section, we examine the PCA-BNN scenario. To address the curse of dimensionality, we applied Principal Component Analysis (PCA) and selected the first 10 principal components as the input features for all models. The neural network architecture consisted of two hidden layers with 10 and 15 neurons, respectively, employing sigmoid activation functions. For the Bayesian models, we assigned independent Gaussian priors over the weights and the variance parameter as $p(\mathbf{W},S)=\mathcal{N}(\mathbf{W};\mathbf{0},100\mathbf{I})\mathcal{N}(S;0,100)$. We compared our proposed method (VBNET-SVAR) against the fixed-variance BNN (VBNET-FIXED), a Standard Neural Network (NNET), and a Generalized Additive Model (GAM).
Results and Discussion: Table 1 summarizes the quantitative results averaged over 10 independent runs. We report the Mean Squared Prediction Error (MSPE), Average Prediction Interval Width (PI Width), and Coverage Probability (CP), along with their Standard Errors (SE).
Model MSPE (SE) Avg PI Width (SE) Coverage Prob. (SE)
VBNET-SVAR 0.7891 (0.0814) 4.5022 (0.0699) 0.9800 (0.0102)
VBNET-FIXED 1.4006 (0.1784) 2.8242 (0.0411) 0.8000 (0.0298)
GAM 1.3396 (0.1839) - -
NNET 0.9177 (0.1274) - -
Table 1: Performance comparison on the Riboflavin dataset (PCA-BNN scenario) over 10 independent runs. Values represent Mean (Standard Error).
As shown in Table 1, VBNET-SVAR achieves the lowest prediction error (MSPE = 0.7891), significantly outperforming VBNET-FIXED (MSPE = 1.4006) and the frequentist baselines. Figure 4 illustrates the predicted values and 95% prediction intervals for four random test samples. The results highlight a key advantage of the proposed method:
  • Uncertainty Quantification: VBNET-FIXED produces overly narrow prediction intervals (Avg Width $\approx$ 2.82), leading to a low coverage probability of 0.80. This suggests the model is overconfident and underestimates the noise introduced by dimensionality reduction.
  • Robustness: In contrast, VBNET-SVAR learns the posterior distribution of the variance, adapting the intervals (Avg Width $\approx$ 4.50) to fully encompass the epistemic uncertainty. This results in a coverage probability of 0.98, demonstrating that the model provides safer and more reliable predictions in this high-uncertainty regime.
Figure 5 further confirms these findings, showing the distribution of MSPE, interval width, and coverage probability across all 10 runs.
Figure 6 illustrates the approximate posterior distribution of the variance parameter for different test samples. By integrating over this distribution during prediction, VBNET-SVAR yields the robust and well-calibrated intervals observed in the results.
Figure 4: Prediction results for samples 1-4 on the Riboflavin dataset (PCA-BNN scenario). The blue dashed lines represent the 95% prediction intervals of the proposed VBNET-SVAR model, which successfully cover the target values.
Figure 5: Performance comparison for the PCA-BNN scenario over 10 independent runs. Left: MSPE (lower is better); Middle: Prediction Interval Width; Right: Coverage Probability (dashed line represents the nominal 95% level).
Figure 6: The approximate Posterior distribution of the likelihood variance parameter ($\sqrt{g(S)}$) for different test samples in the PCA-BNN scenario.
6.3 Riboflavin dataset: Dropout-BNN Scenario
In the second scenario, we utilized the full feature space ($p=4088$) without dimensionality reduction. To handle the high dimensionality, we employed a Bayesian Dropout approach imposed by a spike-and-slab prior. We compared VBNET-SVAR against VBNET-FIXED, a Neural Network with Dropout (DropoutNN), and the sparse-GAM model (GAMSEL) [7]. The network architecture remained two hidden layers with 128 and 64 neurons, and results were averaged over 10 independent runs.
Results and Discussion: Table 2 summarizes the performance metrics. Consistent with the PCA scenario, our proposed method demonstrates superior performance.
Model MSPE (SE) Avg PI Width (SE) Coverage Prob. (SE)
VBNET-SVAR 0.3077 (0.0614) 3.8370 (0.0378) 1.0000 (0.0000)
VBNET-FIXED 0.3607 (0.0603) 1.3412 (0.0216) 0.7200 (0.0490)
GAMSEL 0.4537 (0.0977) - -
NN-Dropout 0.4893 (0.1150) - -
Table 2: Performance comparison on the Riboflavin dataset (Dropout-BNN scenario) over 10 independent runs. Values represent Mean (Standard Error).
The results in Table 2 demonstrate the significant advantage of modeling variance uncertainty in high-dimensional settings:
  1. 1.
    Prediction Accuracy: VBNET-SVAR achieves the best predictive performance with an MSPE of 0.3077, outperforming both the fixed-variance BNN (0.3607) and the frequentist baselines (GAMSEL and NN-Dropout).
  2. 2.
    Uncertainty and Robustness: A striking difference is observed in uncertainty quantification. VBNET-FIXED exhibits a coverage probability of only 0.72, which is far below the nominal 95% level. This indicates severe overconfidence, where the model underestimates the noise (Avg Width $\approx$ 1.34) and fails to capture the true targets. In contrast, VBNET-SVAR adapts to the high epistemic uncertainty inherent in this $p\gg n$ regime. By learning a wider variance, it produces conservative prediction intervals (Avg Width $\approx$ 3.84) that achieve 100% coverage, ensuring that the model is robust and reliable for decision-making.
Figure 7 visualizes these predictions for four random samples. Figure 8 compares the distribution of metrics across runs. The boxplots confirm that VBNET-SVAR consistently maintains lower error and superior coverage compared to the baselines. Additionally, Figure 9 displays the posterior density of the variance parameter ($\sqrt{g(S)}$) learned by VBNET-SVAR, confirming that the method successfully infers the uncertainty associated with the observation noise.
Figure 7: Prediction results for samples 1-4 on the Riboflavin dataset (Dropout-BNN scenario). The blue dashed lines (VBNET-SVAR) provide robust 95% prediction intervals that fully cover the data, while green dashed lines (VBNET-FIXED) are overly narrow.
Figure 8: Performance metrics for the Dropout-BNN scenario over 10 independent runs. Left: MSPE; Middle: PI Width; Right: Coverage Probability. The red dashed line represents the 95% nominal coverage.
Figure 9: The approximate Posterior distribution of the likelihood variance parameter ($\sqrt{g(S)}$) for different test samples in the Dropout-BNN scenario.